mathslover (mathslover):

A triangle and a parallelogram have the same base and the same area. If the sides of the triangle are 15 cm, 14 cm and 13 cm and the parallelogram stands on the base 15 cm, find the height of parallelogram.

hartnn (hartnn):

where are u stuck?

mathslover (mathslover):

I am not stuck but

hartnn (hartnn):

could u find area of triangle?

mathslover (mathslover):

This is a 9th class question and I assume that height will not be in sqrt

mathslover (mathslover):

Area of ABC = 26 sqrt{66} cm^ 2

mathslover (mathslover):

is that right?

hartnn (hartnn):

why am i getting 84

mathslover (mathslover):

\[\large{\sqrt{26(26-15)(26-14)(26-13)}}\]

mathslover (mathslover):

Oh I got the same earlier I put s = 21 but this time I put s = 26

hartnn (hartnn):

semi perimeter 26 ? 21...

hartnn (hartnn):

yes.

mathslover (mathslover):

Oh calculation mistake... sorry

mathslover (mathslover):

I just did that silly mistake ... oops

mathslover (mathslover):

Though Thanks For Helping...

hartnn (hartnn):

np, be careful next time :)

mathslover (mathslover):

yes... thanks

mathslover (mathslover):

height = 84/15 ....

mathslover (mathslover):

or 5.6 cm = h ... correct? ok thanks

hartnn (hartnn):

yes its 5.6

mathslover (mathslover):

k thanks

mathslover (mathslover):

84 = h * 15